How many moles of water are produced when you start with 12.5 moles of oxygen? *
1
Expert's answer
2020-04-28T12:19:56-0400

2H2+O2=2H2O

25H2+12,5O2=25H2O

n(O2)=12,5 moles.

n(H2O)=12,5*2=25 moles